OpenStudy (precal):

Calculus-Tangent Line For what value(s) of x will be the slope of the tangent line to the graph of h(x)=4 times square root (x) be -2? Find the equation of the line tangent to h(x) at this/these x-values. Show your work.

OpenStudy (precal):

\[h(x)=4\sqrt{x}\]

OpenStudy (precal):

I took the derivative of h(x) and set it equal to -2 and found x=1

OpenStudy (precal):

\[h'(x)=\frac{ 2 }{ \sqrt{x} }\]

OpenStudy (precal):

I found h'(1)=2 my point is (1,4) My tangent line is y-4=2(x-1)

OpenStudy (precal):

When I went to check my work, I am not seeing the slope of the tangent line -2 so what am I overlooking?

OpenStudy (mathmale):

Let's take another stab at finding the derivative of\[h(x)=4\sqrt{x}.\]I would first write this function as\[h(x)=4\sqrt{x}=4x ^{1/2}=4(1/2)x ^{(1/2)-1}.\]

OpenStudy (mathmale):

Please finish finding the derivative, and then compare your result with your previous one.

OpenStudy (precal):

\[h'(x)=2x ^{-1/2}\]

OpenStudy (precal):

I thought I did finish finding the derivative above about 3 steps back

OpenStudy (precal):

\[h'(x)=\frac{ 2 }{ \sqrt{x}}\]

OpenStudy (mathmale):

Yes, which is precisely the same as what you got, except for format. The problem statement says that this slope is equal to -2. Mind solving for x again? Would like to see your work.

OpenStudy (precal):

no I don't mind. Please give me a moment to type it in

OpenStudy (precal):

\[\frac{ 2 }{ \sqrt{x} }=-2\]

OpenStudy (precal):

I cross multiplied

OpenStudy (mathmale):

Is there a solution?

OpenStudy (precal):

\[-2\sqrt{x}=2\] divided both sides by -2 \[\sqrt{x}=-1\] squared both sides \[x=1\]

OpenStudy (mathmale):

Sticky. Supposing that you now substitute x=1 into \[\frac{ 2 }{ \sqrt{x} }=-2\]...would the result be true or false?

OpenStudy (precal):

False, ok they got me..........

OpenStudy (precal):

I also, know the parent function is the square root function which is increasing

OpenStudy (mathmale):

It's a little hard to get 2/Sqrt(1) to equal -2, isn't it? Might be worth actually graphing h(x)=4Sqrt(x).

OpenStudy (precal):

yes actually I did graph it and realized that the slope could not be -2 anywhere. I just did not think the question was meant to trick me. Thanks :)

OpenStudy (mathmale):

I think this graph will immediately tell you why the slope of this function h)x) is not going to be -2. You're welcome!

OpenStudy (precal):

I thought I had done something wrong.

OpenStudy (precal):

Another great reason to learn your parent functions.....

OpenStudy (mathmale):

That's always a problem for us bright people, isn't it? We hate to be wrong. Also a good idea to be thoroughly familiar with the shapes of the graphs of common functions, including \[\sqrt{x}, x,x^2, \sin x, \cos x, and so on.\]

OpenStudy (precal):

yes thanks once again. :)

OpenStudy (mathmale):

You're welcome! Bye.